qveentat16:

0.007592 kg=_____mg 1kg = 1000g 1g = 1000mg A) 759.2 B) 7592 C) 75920 D) 759200

justjm:

Use dimensional analysis \(0.007692\ kg\ \cdot\frac{1000\ g}{1\ kg}\cdot\frac{1000\ mg}{1\ g}\) Notice and appreciate how the units at the bottom cancel out, which is the trick to doing these conversions (:

justjm:

@qveentat16 you got that?